For the given functions f and g , find the indicated composition.

f(x) = 3x + 13, g(x) = 4x - 1
(f∘g)(x)

Answer by ewatrrr(24785)


Hi
f(x) = 3x + 13, g(x) = 4x - 1
(f∘g)(x) = 3(4x-1) + 13 = 12x + 10
Wish You the Best in your Studies.